OpenStudy (anonymous):

MEDALS AWARDED! Find the exact real value of arcsin(1/2)

OpenStudy (anonymous):

\[\alpha=\arcsin(1/2)->\sin(\alpha)=1/2->\alpha =\pi/6\]